Would I plug in the numbers below in place of the (w) to solve? How? Given the function H(w)=5w +8, find each of the following.
H(9)=
H(-10)=
H(6.8)=

Answer by Alan3354(69443)

Given the function H(w)=5w +8, find each of the following.
H(9)=
Sub 9 for w:
H(9) = 5*9 + 8 = 53
--------
H(-10)=
Sub -10 for w
-----------
H(6.8)=
Sub 6.8 for w
-------------
PS It's not solving, just evaluating.